OpenStudy (ammarah):

Solve the equations: 1. 2x^ +3x-2=0 2. 2x^-3x-9=0 3. 4x^-8x+3=0 4. 9x^-4=0

OpenStudy (anonymous):

Youre missing the exponents...

OpenStudy (ammarah):

@mathstudent55

OpenStudy (ammarah):

^ means 2

OpenStudy (ammarah):

squared....

OpenStudy (mathstudent55):

Please write x^2 to mean an exponent of 2 like everyone does. Otherwise, it gets confusing.

OpenStudy (mathstudent55):

Let's start with the first one.

OpenStudy (ammarah):

ohk im sorry

OpenStudy (mathstudent55):

\(2x^2 +3x-2=0\) We need to factor the above. You can try the guess method or the ac method.

OpenStudy (ammarah):

ok ill tell u what i get for an answer

OpenStudy (mathstudent55):

Since 2x^2 can only be the product of 2x and x, the guess method here would not be too long.

OpenStudy (ammarah):

i got x=-2or 1/2

OpenStudy (mathstudent55):

\(2x^2 +3x-2=0\) \( = (2x~~~~~~)(x~~~~~~) \) The numbers that need to go in the parentheses have to multiply to -2. There are only two possibilities, -2 and 1 or -1 and 2. Now we try to see which two numbers and in which position they will work.

OpenStudy (mathstudent55):

You are correct.

OpenStudy (ammarah):

for the second one i got x=-3/2 or 3

OpenStudy (mathstudent55):

\( (2x - 1)(x + 2) \) x = 1/2, or x = -2

OpenStudy (mathstudent55):

Second one is correct.

OpenStudy (ammarah):

ok i get number 3 i need help with 4

OpenStudy (mathstudent55):

4. is the type of factoring that is called the "difference of two squares." Here is the pattern for the factorization of the difference of two squares: \( a^2 - b^2 = (a + b)(a - b) \) Example: Factor \(x^2 - 25\) Solution: \(x^2 - 25 = x^2 - 5^2 = (x + 5)(x - 5)\)

OpenStudy (mathstudent55):

Now look at your problem 4. Is this the difference of two squares? Is the first term, 9x^2, the square of something? Yes, it's the square of 3x. Is the second term, 4, the square of something? Yes, it's the square of 2. That means the problem can be written as \((3x)^2 - 2^2 \) Now you can see that it factors as follows: \(9x^2 - 4 = (3x)^2 - 2^2 =(3x + 2)(3x - 2) \)

OpenStudy (ammarah):

okay i dont know how to do 6

OpenStudy (mathstudent55):

I don't either because you only listed 4 problems above.

OpenStudy (ammarah):

oh oops 9x^2 +11x+18=-10x+8

OpenStudy (mathstudent55):

First, you need to bring all terms to the left side, combione like terms, and make the expression on the left equal zero.

OpenStudy (mathstudent55):

\(9x^2 +11x+18=-10x+8\) \(9x^2 +11x + 10x +18 - 8 = 0\) \(9x^2 + 21x + 10 = 0\)